# Private Fine-tuning of Large Language Models with Zeroth-order Optimization

### v1

@@ 3 OUR METHOD: DP-ZO
Table 1: Main results with 1000 training samples for each dataset. OPT-13B model with LoRA fine-tuning. DP-ZO (G) is DP-ZO instantiated with the Gaussian mechanism. $\delta=10^{-5}$. The $\varepsilon=\infty$ by ZO is $84.73$ for SQuAD, $30.78$ for DROP, and $92.09$ for SST2. The $\varepsilon=0$ baseline, i.e., directly doing model evaluation without training, is $46.23$ for SQuAD, $14.64$ for DROP, and $58.83$ for SST2.
@@
DP-ZO(G) $79.22_{0.64}$ $81.35_{0.38}$ $82.85_{0.31}$ $28.27_{0.40}$ $29.45_{0.40}$ $30.34_{0.54}$ $85.07_{3.31}$ $89.68_{0.82}$ $92.00_{0.94}$
Table 2: DP-ZO (Gaussian) and DP-SGD with full parameter and LoRA fine-tuning on SQuAD with 1000 training samples across different model sizes. $(1,10^{-5})$-DP. ‘-’ means the approach did not scale with straightforward implementation;  Section 5 details the additional engineering required to scale DP-SGD to larger models. ‘- -’ for DP-ZO means the results are omitted due to limited computational resources.
@@
DP-ZO-LoRA (Gaussian) $74.41_{0.81}$ $79.30_{0.93}$ $79.53_{0.65}$ $81.35_{0.38}$ $82.03_{0.90}$ $83.87_{0.99}$
DP-SGD-LoRA $75.39_{0.33}$ $79.42_{0.57}$ $79.53_{0.52}$ $82.14_{0.18}$
DP-ZO-Full (Gaussian) $72.84_{1.03}$ $77.25_{0.27}$ $79.06_{0.67}$ $82.16_{0.41}$
DP-SGD-Full $75.50_{0.89}$ $79.81_{0.64}$
4 EVALUATION
We first overview our experimental setup in Section 4.1 and then evaluate the performance of DP-ZO in Section 4.2. We find that DP-ZO provides a competitive privacy-utility tradeoff for conservative privacy budgets across multiple datasets and model scales. One noteworthy result is that DP-ZO exhibits just $1.86\%$ accuracy degradation due to privacy at $\varepsilon=1$ when fine-tuning OPT-66B on 1000 training samples from SQuAD. We also compare DP-ZO to DP-SGD in Section 4.2. In Section 4.3 we first characterize DP-ZO under different few-shot settings. We then show that DP-ZO is robust to different mechanisms for noise addition. Specifically, DP-ZO with the Laplace mechanism is the first method to achieve a non-trivial privacy-utility trade-off under pure $\varepsilon$-DP. We also discuss the choice of batch size for DP-ZO.
4.1 Experimental Setup
When we report results, we report the metric of interest (F1 score or accuracy) and standard deviation averaged across 5 independent runs with different random seeds. We detail the full hyperparameter searches in Appendix D.
Datasets.
We consider three different benchmark NLP tasks: SQuAD (Rajpurkar et al., 2016) and DROP (Dua et al., 2019) for text generation, and SST2 (Socher et al., 2013) for text classification. Although all three datasets have very different dataset sizes, we consider the few-shot setting for all three datasets where we sample $1000$ points for each dataset. Fine-tuning LLMs with $n$ samples on the order of $n=1000$ is a standard setting widely studied in the NLP community (Gao et al., 2021; Malladi et al., 2023) because we are generally interested in the few-shot abilities of LLMs (Brown et al., 2020). This represents a departure from prior works that privately finetune LLMs; Yu et al. (2022); Li et al. (2022b); Yu et al. (2021b) use the entire training dataset of SST2 that has about 65,000 examples. It is well known that the privacy-utility tradeoff improves greatly with more data (Tramèr and Boneh, 2021). It is straightforward to see that our setting with datasets of the size $n=1000$ with $\delta=10^{-5}$ is simultaneously more challenging and more aligned with real-world usecases than previous works in DP finetuning of LLMs. Despite the increased difficulty of our few-shot setting as compared to prior work, our results validate that DP-ZO realizes a strong privacy-utility tradeoff.
Models.
We use models from the OPT (Zhang et al., 2022) family of decoder-only transformers. We present our main results (Table 1) using a pretrained OPT-13B model that is finetuned with LoRA (Hu et al., 2022); that is, we update $<1\%$ of the total parameters. We include a range of ablation studies, including varying the model size and amount of parameters to be updated after we present main results.
Privacy Budgets.
We consider various privacy levels with $\varepsilon=[0.5,1,4]$ and fix $\delta=10^{-5}$ for $(\varepsilon,\delta)$-DP. We include the zero-shot $\varepsilon=0$ baseline does not incur any privacy loss because we evaluate the pretrained model directly without finetuning on private data. We also include the non-private $\varepsilon=\infty$ baseline that is trained without any DP guarantee. That is, we iterate over the shuffled dataset instead of doing Poisson sampling (replacing line 4), do not clip the step size (skipping line 11) and set $\sigma=0$ (in line 12). We make these modifications between Poisson sampling and clipping are known to degrade performance, and we want to compare to the strongest possible nonprivate baseline.
4.2 Main Results
DP-ZO Provides a Strong Privacy-utility Tradeoff for Conservative Privacy Budgets.
As shown in Table 1, across all three tasks and all $\varepsilon$s, DP-ZO significantly improves upon the $\varepsilon=0$ baseline, and only slightly degrades the performance compared to the non-private baseline. For SQuAD, even at $\varepsilon=0.5$, DP-ZO can still achieve $78.10\%$, that significantly outperforms $\varepsilon=0$ baseline ($46.23\%$). The gap between $\varepsilon=0.5$ and $\varepsilon=\infty$ is about $5.5\%$. By increasing $\varepsilon$ from $0.5$ to $4$, this gap can be further reduced to $2\%$. For DROP and SST2, DP-ZO (Gaussian) achieves comparable performance as the non-private baseline at $\varepsilon=4$.
DP-ZO Scales to Large Models.
In Table 2 we show that DP-ZO continues improving as the model size increases from 1.3B to 66B. Due to space constraints, we provide the non-private ($\varepsilon=\infty$) performance of all models and methods in Appendix E. Table 2 shows an promising insight: as the model size and nonprivate performance increase, the gap in performance between private and nonprivate models shrinks. Specifically, the gap for OPT-125M is $4.96\%$ ($48.7\%$ at $\varepsilon=\infty$ reduced to $43.74\%$ under $\varepsilon=1$; omitted from Table 2 for poor performance). But this gap shrinks to just $1.86\%$ for OPT-66B, where the private performance at $\varepsilon=1$ is $83.87\%$ compared to $85.73\%$ non-privately. Our findings therefore suggest that DP-ZO scales to large models not only because it is compatible with existing pipeline without much additional engineering effort but also because the utility drop due to privacy is smaller as the model size increases.
Comparison with DP-SGD.
@@ 4.2 Main Results
We compare DP-ZO to differentially private stochastic gradient descent (DP-SGD) (Abadi et al., 2016) which has been applied to fine-tune LLMs with full parameter fine-tuning (Li et al., 2022b) and with LoRA (Yu et al., 2022; He et al., 2023). Recall that DP-ZO is compatible out-of-the-box with mixed precision training and GPU parallelism, enabling us to fine-tune OPT-66B. As we show in  Section 5, it is significantly more challenging to integrate DP-SGD with these techniques, and furthermore, DP-SGD requires more memory than DP-ZO to store activations and compute per-sample gradients. As a direct result, DP-SGD cannot directly scale past 2.7B with full fine-tuning or 13B with LoRA without additional implementations desired for multiple GPUs, while DP-ZO can scale seamlessly to larger models. In Table 2 we present comparisons between DP-ZO and DP-SGD with full parameter finetuning and LoRA. The best performance by DP-ZO is  83.87% by OPT-66B finetuned with LoRA. This is $\approx 2\%$ better than the best performance of DP-SGD in Table 2 that is $82.14\%$ by OPT-13B with LoRA. Note that when compared to the same model size, the performance of DP-ZO is $1\%$ worse than DP-SGD with LoRA and $2-3\%$ worse than DP-SGD with full parameter funetuning. How to close the gap between DP-ZO and DP-SGD for the same model size is an interesting direction and we leave it as future work.
@@ 4.3 Analysis
We vary the amount of training data we sample, the choice of DP mechanism in DP-ZO, and the batch size.
@@
$\varepsilon=1$ $73.86_{0.76}$ $77.43_{0.51}$ $81.35_{0.38}$ $82.35_{0.48}$
$\varepsilon=\infty$ $84.03$ $84.66$ $84.73$ $84.75$
Characterizing Differentially Private Few-Shot Learning.
Although it is known that private learning requires more data than non-private learning (Bassily et al., 2014), prior work has not characterized this improvement for fine-tuning language models. In Fig. 1 and Table 3 we vary the number of training samples $n$ around the $n=1000$ setting in the main results while keeping $\delta=10^{-5}$ fixed for all choices of $n$ such that the largest $n=5000$ still satisfies $\delta\ll 1/n$.
Table 3 shows that DP-ZO can achieve nontrivial performance in few-shot settings under conservative privacy guarantees. Furthermore, we find that while increasing the amount of training data by 10×10$\times$10 × barely increases non-private performance, it increases private performance by $\approx 5\%$. While non-private few-shot learning can succeed by just memorizing the training data, Fig. 1 indicates that the convergence rate for different shots for private few-shot learning is different. With the proliferation of pretrained models, we anticipate that privately fine-tuning downstream tasks in the few-shot setting will be more aligned with real-world use cases.
Table 4: Pure $\varepsilon$-DP by DP-ZO (Laplace), SQuAD with 1000 training samples. OPT-13B with LoRA fine-tuning. The $\varepsilon=\infty$ performance with ZO is $84.73\%$ and the $\varepsilon=0$ baseline is $46.23\%$.
$\varepsilon$ $\varepsilon=4$ $\varepsilon=10$ $\varepsilon=15$
DP-ZO (Laplace) $64.50_{1.04}$ $73.69_{1.25}$ $76.99_{0.76}$
DP-ZO with Pure $\varepsilon$-DP.
To the best of our knowledge, DP-ZO (Laplace) is the first method that achieves a non-trivial privacy-utility tradeoff under pure $\varepsilon$-DP. In Table 4 we find that DP-ZO (Laplace) can significantly improve upon $\varepsilon=0$, albeit for a large privacy budget. Given a budget $\varepsilon=10$, which some prior work has considered reasonable (Ponomareva et al., 2023), DP-ZO (Laplace) can obtain $73.69\%$ on SQuAD. As we showed in Table 3, the few-shot setting poses a unique challenge for obtaining strong performance under conservative privacy budgets. Acquiring more data enhances privacy amplification and reduces the amount of noise we need to add to achieve a target $\varepsilon$-DP guarantee (Figure 2). In particular, in Table 5 we find that increasing the number of training examples from $1000$ to $5000$ improves performance at $\varepsilon=4$ from $64.50\%$ to $78.65\%$, although the improvement of non-private performance at $\varepsilon=\infty$ by increasing training samples from $1000$ to $5000$ is insignificant. DP-ZO (Laplace) presents a new direction for obtaining privacy guarantees without failure probabilities.
Table 5: Pure $\varepsilon$-DP by DP-ZO (Laplace) at $\varepsilon=4$, SQuAD with different training samples. OPT-13B with LoRA fine-tuning. The $\varepsilon=\infty$ by ZO is $84.73\%$ and $84.75\%$ for 1000 and 5000 samples respectively. The $\varepsilon=0$ baseline is $46.2\%$.
@@
DP-ZO (Laplace) $64.50_{1.04}$ $78.65_{1.37}$
Figure 2: The added noise scale $\sigma$ for different number of samples to satisfy $(4,0)$-DP decreases with data size.
@@
$\varepsilon$ $\varepsilon=0.35$ $\varepsilon=0.5$ $\varepsilon=1$ $\varepsilon=4$
DP-ZO (G) $76.90_{0.23}$ $79.22_{0.64}$ $81.35_{0.38}$ $82.85_{0.31}$
DP-ZO (L) $73.69_{1.25}$ $76.76_{0.74}$ $80.10_{0.85}$ $82.70_{0.39}$
Different Noise Mechanisms for $(\varepsilon,\delta)$-DP.
We now relax the privacy guarantee provided by the Laplace mechanism to approximate $(\varepsilon,\delta)$-DP. For example, the privacy parameter setting given by basic composition as $(10,0)$-DP in Table 4 can also be tracked via composition of private random variable for privacy curve of a $\varepsilon$-DP algorithm and costs $(0.35,10^{-5})$-DP. In Table 6, we compare DP-ZO instantiated with the Laplace and Gaussian mechanisms. DP-ZO (Gaussian) outperforms DP-ZO (Laplace) for strict privacy budgets such as $\varepsilon=0.35$ because it enjoys tighter accounting (Gopi et al., 2021) and lower variance (Dwork and Roth, 2014). These advantages are less significant for larger privacy budgets; for $\varepsilon=4$, both DP-ZO (Gaussian) and DP-ZO (Laplace) exceed $82\%$.
Ablation Study of Batch Size for DP-ZO.
The tradeoff between decreasing the batch size to benefit from privacy amplification by subsampling (Balle et al., 2018) and increasing batch size to minimize the per-example added noise (De et al., 2022; Ponomareva et al., 2023) has trended towards training with large batches (Li et al., 2022b) or even doing full-batch gradient descent (Panda et al., 2022). However, these analyses are specific to SGD, and the optimal batch size for DP-ZO is of course an open question.
Figure 3: (Smoothed) training loss. Larger batch sizes have a slightly better convergence rate and the convergence rates for batch size of 16, 32, 64 are similar.
In Fig. 3 and Table 7 we ablate the batch size $B$ in DP-ZO (Gaussian) around the batch size $B=16$ we use for our main results. Somewhat surprisingly, we find that increasing the batch size does not strictly improve performance for $(1,10^{-5})$-DP, even though it has been shown to improve performance in the non-private setting (Liu et al., 2018).
Table 7: Different batch size $B$ on SQuAD 1000 training samples. OPT-13B model. $\varepsilon=1$. 10,000 steps.
$B$ 4 8 16 32 64
$\sigma$ $1.68$ $3.11$ $6.08$ $12.07$ $24.09$
F1 $79.94$ $80.95$ $81.35$ $81.63$ $81.47$
Std. $0.29$ $0.63$ $0.38$ $0.62$ $0.81$
One possible reasoning for this is the nature of zeroth-order optimization. Under DP-SGD, increasing the batch size decreases the variance of the gradient estimator and therefore decreases the number of iterations required for convergence (Bassily et al., 2020), which in turn reduces the amount of noise that must be added at each iteration of DP-SGD to ensure privacy (Kairouz et al., 2015) and counterbalances the loss of amplification. Note that this does not reduce the actual amount of computation done, because DP-SGD typically requires gradient accumulation to train large models with large batch sizes (De et al., 2022). By contrast, because SPSA inherently has higher variance than SGD (Malladi et al., 2023), the variance reduction by large batch size in DP-ZO may not be as significant as on DP-SGD. An interesting direction for future work is designing a batch size schedule around the intricacies of SPSA. In Appendix D we did an initial study to systematically evaluate the interplay between batch size and training iterations for DP-ZO.

### v2

@@ 4.2 Main Results
Table 1: Main results with 1000 training samples for each dataset. OPT-13B model with LoRA fine-tuning. DP-ZO (G) is DP-ZO instantiated with the Gaussian mechanism. $\delta=10^{-5}$. The $\varepsilon=\infty$ by ZO is $86.85$ for SQuAD, $33.22$ for DROP, and $93.69$ for SST2. The $\varepsilon=0$ baseline, i.e., directly doing model evaluation without training, is $46.23$ for SQuAD, $14.64$ for DROP, and $58.83$ for SST2. The results of DP-SGD on DROP are omitted because fine-tuning OPT-13B on the DROP dataset by LoRA will cause the out of memory issue on a single A100 GPU even in the non-private setting.
@@
DP-ZO(G) $80.10_{0.63}$ $82.28_{0.84}$ $83.87_{0.50}$ $28.39_{0.82}$ $30.30_{0.51}$ $31.99_{0.51}$ $85.41_{2.91}$ $91.19_{0.90}$ $92.59_{0.30}$
DP-SGD $79.85_{0.89}$ $82.14_{0.18}$ $83.05_{0.51}$- -- -- $64.33_{6.47}$ $90.25_{0.78}$ $92.06_{0.52}$
DP-ZO scales to large models. In Table 2 we show that DP-ZO continues improving as the model size increases from 1.3B to 66B. Due to space constraints, we provide the non-private ($\epsilon=\infty$) performance of all models and methods in Appendix E. Table 2 shows an promising insight: as the model size and nonprivate performance increase, the gap in performance between private and nonprivate models shrinks. Specifically, the gap for OPT-1.3B is $5.68\%$ ($80.97\%$ at $\varepsilon=\infty$ reduced to $75.29\%$ under $\varepsilon=1$). But this gap shrinks to just $3.37\%$ for OPT-66B, where the private performance at $\epsilon=1$ is $84.12\%$ compared to $87.49\%$ non-privately. Our findings suggest that DP-ZO scales to large models not only because it is compatible with existing pipeline without much additional engineering effort but also because the utility drop due to privacy is smaller as the model size increases.
Table 2: DP-ZO (Gaussian) and DP-SGD with full parameter and LoRA fine-tuning on SQuAD with 1000 training samples across different model sizes. $(1,10^{-5})$-DP. ‘--’ means the approach did not scale with straightforward implementation;  Section 5 details the additional engineering required to scale DP-SGD to larger models. ‘--- -’ for DP-ZO means the results are omitted due to limited computational resources. Due to limited computing resources, this table does not include the standard deviation for OPT-66B model.
@@
DP-ZO-LoRA (Gaussian) $75.29_{0.90}$ $80.34_{1.14}$ $81.34_{1.04}$ $82.28_{0.84}$ $82.48_{0.83}$ $84.12_{1.01}$
DP-SGD-LoRA $75.39_{0.33}$ $79.42_{0.57}$ $79.53_{0.52}$ $82.14_{0.18}$- --
DP-ZO-Full (Gaussian) $72.84_{1.03}$ $77.25_{0.27}$ $79.06_{0.67}$ $82.16_{0.41}$-- - --- -
DP-SGD-Full $75.50_{0.89}$ $79.81_{0.64}$- -- -- --
@@
Comparison with DP-SGD. We compare DP-ZO to differentially private stochastic gradient descent (DP-SGD) (Abadi et al., 2016) which has been applied to fine-tune LLMs with full parameter fine-tuning (Li et al., 2022b) and with LoRA (Yu et al., 2022; He et al., 2023). Recall that DP-ZO is compatible out-of-the-box with mixed precision training and GPU parallelism, enabling us to fine-tune OPT-66B. As we discuss in  Section 5, it is significantly more challenging to integrate DP-SGD with these techniques, and furthermore, DP-SGD requires more memory than DP-ZO to store activatiosn and compute per-sample gradients (see Section 4.3). As a direct result, DP-SGD cannot directly scale past 2.7B with full fine-tuning or 13B with LoRA without additional implementation effort for multi-GPU training, while DP-ZO can scale seamlessly to larger models. In Table 2 we present comparisons between DP-ZO and DP-SGD with full parameter finetuning and LoRA. With the same model size, DP-ZO achevies comparable performance as DP-SGD as by LoRA finetuning, i.e., both DP-ZO and DP-SGD achieves $82\%$ on OPT-13B models. The best performance by DP-ZO is $84.12\%$ by OPT-66B finetuned with LoRA. This is $\approx 2\%$ better than the best performance of DP-SGD in Table 2 that is $82.14\%$ by OPT-13B with LoRA.
DP-ZO with pure $\varepsilon$-DP. To the best of our knowledge, DP-ZO (Laplace) is the first method that achieves a non-trivial privacy-utility trade-off under pure $\varepsilon$-DP under a conservative privacy budget like $\varepsilon=4$ on large language models.
In Table 3, DP-ZO (Laplace) can significantly improve upon $\varepsilon=0$. Given a budget $\varepsilon=4$, which some prior work has considered reasonable (Ponomareva et al., 2023), DP-ZO (Laplace) can obtain $73.52\%$ on SQuAD. When increasing $\varepsilon=4$ to $\varepsilon=15$, DP-ZO (Laplace) can obtain $78.82\%$ on SQuAD. Note that the $l_{1}$ sensitivity required for Laplace mechanism makes it hard to DP-SGD to achieve comparable performance as DP-ZO because the gradients in DP-SGD have high dimension. Table 3 shows that DP-SGD with $l_{1}$ norm clipping and Laplace noise only achieves $47.25\%$ for reasonable privacy budgets with $\varepsilon$ ranging from $4$ to $15$, that is only marginal improvement upon the zero-shot performance. Even when relaxing the privacy budget to near-vacuous guarantees such as $\varepsilon=10450$, DP-SGD (Laplace) still achieves worse performance compared to DP-ZO due to the Laplace noise added in high-dimension gradients.
Table 3: Pure $\varepsilon$-DP by DP-ZO (Laplace), SQuAD with 1000 training samples. OPT-13B with LoRA fine-tuning. The $\varepsilon=0$ baseline is $46.23\%$.
$\varepsilon$ $\varepsilon=4$ $\varepsilon=10$ $\varepsilon=15$ $\varepsilon=10450$
DP-ZO (Laplace) $73.52_{1.04}$ $76.75_{1.39}$ $78.82_{1.57}$ $81.02_{1.24}$
DP-SGD (Laplace) $47.25_{0.79}$ $47.27_{0.95}$ $47.36_{1.02}$ $76.50_{0.89}$
@@ 4.3 Analysis
In this section, we provide ablation studies on the feasibility of DP-ZO across different model architectures, empirical privacy analysis to measure how private is DP-ZO, memory efficiency of DP-ZO, the amount of training data that we sample in DP-ZO, and the choice of DP mechanism in DP-ZO.
DP-ZO provides a strong privacy-utility trade-off across different model architectures. Table 1 and Table 2 show that DP-ZO achieves the comparable performance as DP-SGD on various OPT models sizes. We now run experiments on SQuAD with Mistral-7B-v1 model (Jiang et al., 2023) in Table 4 and include the results of OPT-6.7B and OPT-13B for the ease of comparison. DP-ZO and DP-SGD both achieve comparable performance at $\varepsilon=1$, that are around $89\%$. Moreover, even Mistral-7B-v1 has similar model parameters as OPT-6.7B and much fewer parameters than OPT-13B, DP-ZO achieves better performance by Mistral-7B-v1 than OPT-13B. This indicates that with the development of more advanced models, the power of DP-ZO will be further unlocked.
Table 4: Ablating DP-ZO across different model architectures. The $\varepsilon=0$ baseline for Mistral-7B-v1 is 68.37.
Models OPT-6.7B OPT-13B Mistral-7B-v1
DP-ZO $81.34_{1.04}$ $82.28_{0.84}$ $89.79_{0.41}$
DP-SGD $79.53_{0.52}$ $82.14_{0.18}$ $89.44_{0.41}$
DP-ZO is memory efficient. We omitted several results of DP-SGD in Section 4.2 due to excessive memory consumption of DP-SGD that leads to out-of-memory (OOM) on a single A100 80G GPU. We now provide a more fine-grained memory analysis for a better understanding of DP-ZO and DP-SGD.
The naive implementation of DP-SGD causes additional memory consumption due to the per-example gradient computation. An ongoing line of work (Li et al., 2022b; Yu et al., 2022; He et al., 2023; Bu et al., 2024) has continued improving the scalability of DP-SGD over the past few years. For memory cost comparison, we consider several variants of DP-SGD including DP-SGD-full (Abadi et al., 2016; Li et al., 2022b), DP-SGD-full(ghost) (Li et al., 2022b), DP-SGD-LoRA (Yu et al., 2022), DP-SGD-BitFiT (Bu et al., 2024) and DP-ZO (including full and LoRA) for a fair comparison. [^1] DP-SGD-full, DP-SGD-full(ghost), DP-SGD-LoRA are based on https://github.com/lxuechen/private-transformers. For DP-SGD-BitFit, we use fastDP https://github.com/awslabs/fast-differential-privacy/tree/main/fastDP and follow the guideline by using one line of code [param.requires_grad_(False) for name, param in model.named_parameters() if ’.bias’ not in name].
As discussed in (Li et al., 2022b; De et al., 2022), small batch size will incur sub-optimal performance of DP-SGD and therefore large batch size is preferred, we consider gradient accumulation in memory analysis to enable large batch size. We consider full precision both for DP-SGD and DP-ZO for fair comparison. We present the results of such memory consumption for different sequence lengths on OPT-2.7B in Table 5.
Table 5: Comparison of memory consumption (GB) of DP-ZO and DP-SGD varying differnet sequence length on OPT-2.7B. Full-precision, no gradient check-pointing. Batch size=2, gradient accumulation steps=2. OOM indicates out-of-memory on a single A100 80G GPU.
Methods DP-SGD-full DP-SGD-full(ghost) DP-SGD-LoRA DP-SGD-BitFit DP-ZO-full DP-ZO-LoRA
seq_len=128 $51.3$ $32.9$ $11.4$ $12.2$ $11.6$ $10.3$
seq_len=512 $51.4$ $39.6$ $18.1$ $21.3$ $11.7$ $11.1$
seq_len=2048 OOM OOM OOM OOM $15.6$ $15.6$
Table 5 shows that the naive implementation of DP-SGD incurs around 50GB for sequence length equal to 128 and incurs out-of-memory (OOM) issue when increasing sequence length to 2048. Ghost clipping can help reduce the memory consumption by removing per-sample gradient computation, but still needs to accumulate gradients and consumes memory more than 30GB. Parameter efficient fine-tuning methods including DP-SGD-LoRA, DP-SGD-BitFiT can largely reduce the memory cost for gradient accumulation with fewer parameters in gradients. However, as the input sequence length increases, the activation saved in forward-pass for gradients computation significantly increases and still causes OOM error for sequence length equals to 2048. Note that such long sequence input exists in practical scenarios such as digesting from a long documents and recent foundation models put efforts to support longer context (Gemini-Team et al., 2023). In contrast, DP-ZO does not need to store gradients nor store additional activation for gradients, therefore can reduce the memory consumption to be less than 16GB even when sequence length is 2048. In fact, DP-ZO incurs nearly no additional memory cost than ZO (Malladi et al., 2023).
Figure 5: Memory comparison of DP-ZO and DP-SGD with half-precision and gradient checkpointing. Batch size=1, gradient accumulation steps=2.
We now consider a more restrictive memory setting such as on-device mobile setting, i.e., 8GB as a memory limit (Gim and Ko, 2022; Guo et al., 2024). As discussed in Malladi et al. (2023), gradient checkpointing can help reduce the activation memory consumption. We use gradient checkpointing to reduce the memory consumption of activation and half-precision to reduce the memory consumption of weights and gradients. We measure with the default implementation of activation checkpointing where we checkpoint every block of the model. We consider batch size=1 and accumulate gradients in 2 steps to enable training in large batch size. We present the results of such memory consumption for different sequence length on OPT-2.7B model in Figure 5. While gradient checkpointing largely reduces the memory consumption in DP-SGD, DP-SGD-full is still beyond the 8GB memory limit due to gradient accumulation. Similarly, with gradient checkpointing, DP-SGD-LoRA still exceeds the 8GB limit when increasing sequence to 2048. In contrast, DP-ZO incurs just 6.2GB even when input sequence length is 2048.
How private is DP-ZO and ZO: an empirical privacy analysis. We discussed earlier in Section 2.2 that the single scalar information from Zeroth-order Optimization will leak private information and therefore motivate our design of DP-ZO. We now validate our design by conducting empirical privacy evaluation through membership inference attacks (MIA) (Shokri et al., 2017) to understand the privacy implication of DP-ZO. We use the state-of-the-art privacy auditing empirical privacy evaluation method in Panda et al. (2024b) for empirical privacy evaluation. Similar to Panda et al. (2024b), we construct synthetic canaries by creating one new token for each canary to the vocabulary to increase the information from canaries for better auditing. Following Panda et al. (2024b); Mireshghallah et al. (2022), we use MIA Area under the ROC Curve (AUC-ROC) for empirical privacy evaluation. We report the results for DP-ZO and DP-SGD in Table 6 for different $\varepsilon$s including $[0.5,1,4,10,\infty]$.
As observed in Panda et al. (2024b), such privacy attack is very successful for DP-SGD when no noise added, and DP-SGD can effectively reduce the attack AUC to 54.8 that is closed to random guess at $\varepsilon=0.5$. For DP-ZO, when no noise added, the MIA AUC is 71.4 that is lower than DP-SGD($\varepsilon=\infty$), however much higher than the random guess baseline 50. Interestingly, in this empirical privacy case study, the privacy leakage of ZO is similar to the privacy leakage of DP-SGD at $\varepsilon=10$. DP-ZO is motivated to get a formal privacy guarantee by differential privacy. DP-ZO can reduce the attack AUC to around random guess at $\varepsilon=[0.5,1,4]$. To the best of our knowledge, this is the first experimental result that measures the privacy leakage in ZO and DP-ZO. This result shows the necessity of DP-ZO to redcue MIA cloase to random guess, and also raises an open problem about the inherent privacy property of zeroth-order optimization.
Table 6: Membership Inference Atttack AUC-ROC for DP-ZO and DP-SGD.
$\varepsilon=0.5$ $\varepsilon=1$ $\varepsilon=4$ $\varepsilon=10$ $\varepsilon=\infty$
DP-ZO 53.2 54.5 55.6 58.8 71.4
DP-SGD 54.8 55.0 61.5 71.9 100.0
Characterizing the effect of data size. Although it is known that private learning requires more data than non-private learning (Bassily et al., 2014), prior work has not characterized this improvement for fine-tuning language models. In Table 8 and Figure 6 we first vary the number of training samples $n$ around the $n=1000$ setting in the main results while keeping $\delta=10^{-5}$ fixed for all choices of $n$. Table 8 shows that DP-ZO can achieve nontrivial performance in few-shot settings under conservative privacy guarantees. Furthermore, we find that while increasing the amount of training data by 10×10$\times$10 × barely increases non-private performance, it increases private performance by $\approx 6\%$ ($n=500$ vs. $n=5000$). Similarly, acquiring more data enhances privacy amplification and reduces the amount of noise we need to add to achieve a target $\varepsilon$-DP + guarantee. In Table 8 we find that increasing the number of training examples from $1000$ to $5000$ improves performance at $\varepsilon=4$ from $73.52\%$ to $79.89\%$, although the improvement of non-private performance at $\epsilon=\infty$ by increasing training samples from $1000$ to $5000$ is insignificant.
While non-private few-shot learning can succeed by just memorizing the training data, Figure 6 indicates that the convergence rate for different shots for private few-shot learning is different. With the proliferation of pretrained models, we anticipate that privately fine-tuning downstream tasks in the few-shot setting will be more aligned with real-world use cases.
Besides the few-shot setting, we now scale the size of training data size up to more than 100k samples and compare the performance of DP-SGD and DP-ZO. For the ease of comparison, we follow Li et al. (2022b); Yu et al. (2022) and conduct experiments for DP-ZO and DP-SGD on QNLI with RoBERTa-base with a range of examples at $\varepsilon=3$ and report result in Table 10. Similar to previous observation, Table 10 shows that at a small data regime, i.e., samples=1000, DP-ZO achieves comparable performance as DP-SGD. DP-ZO can also improve its performance within more samples. Compared to DP-SGD, we notice that when increasing the data size, there is a utility gap between DP-ZO and DP-SGD. This limitation shows an open problem that how to improve data efficiency in zeroth-order optimization (Zhang et al., 2024c; Zhao et al., 2024) and DP-ZO, and we leave this data efficiency improvement in ZO as future work.
@@
$\varepsilon=1$ $74.86_{0.74}$ $78.25_{2.38}$ $82.28_{0.84}$ $84.29_{0.92}$
$\varepsilon=\infty$ $86.40$ $86.53$ $86.85$ $86.92$
Table 8: Pure $\varepsilon$-DP by DP-ZO (Laplace) at $\varepsilon=4$, SQuAD with different training samples. OPT-13B with LoRA fine-tuning. The $\varepsilon=\infty$ by ZO is $86.85\%$ and $86.92\%$ for 1000 and 5000 samples respectively. The $\varepsilon=0$ baseline is $46.2\%$.
@@
DP-ZO (Laplace) $73.52_{1.04}$ $79.89_{0.49}$
Table 9: Comparison of DP-ZO and DP-SGD on different samples on QNLI dataset with Roberta-base model.
1000 5000 10000 50000 104743
DP-SGD 73.27 79.44 80.54 84.81 87.40
DP-ZO 76.19 78.66 79.28 79.70 79.85
@@
$\varepsilon$ $\varepsilon=0.5$ $\varepsilon=1$ $\varepsilon=4$
DP-ZO (G) $80.10_{0.63}$ $82.28_{0.84}$ $83.87_{0.50}$
DP-ZO (L) $77.58_{0.81}$ $80.49_{0.63}$ $82.94_{0.69}$
Different noise mechanisms for $(\varepsilon,\delta)$-DP. We now relax the privacy guarantee provided by the Laplace mechanism to approximate $(\varepsilon,\delta)$-DP. In Table 10, we compare DP-ZO instantiated with the Laplace and Gaussian mechanisms. DP-ZO (Gaussian) outperforms DP-ZO (Laplace) for strict privacy budgets such as $\varepsilon=0.5$ because it enjoys tighter accounting (Gopi et al., 2021) and lower variance (Dwork and Roth, 2014). These advantages are less significant for larger privacy budgets; for $\varepsilon=4$, the gap between DP-ZO (Gaussian) and DP-ZO (Laplace) is within $1\%$.
Our ablation study on the DP-ZO with laplace for $\varepsilon$-DP and the comparisons of Laplace and Gaussian mechanisms for $(\varepsilon,\delta)$-DP shows that DP-ZO provides a strong privacy-utility trade-off under different DP mechanisms while DP-SGD suffers from Laplace mechanisms for $(\varepsilon,\delta)$-DP, which opens the new opportunity for the synergy between DP mechanisms and large language models.